OpenStudy (anonymous):

Partial Derivative:

OpenStudy (anonymous):

\[\frac{ \partial f }{ \partial x } e ^{\frac{ x }{ y }}\]

OpenStudy (anonymous):

@abhrag321

OpenStudy (anonymous):

is it \[\frac{ 1 }{ y }e ^{\frac{ x }{ y }}\] ?

OpenStudy (anonymous):

what is f ?

OpenStudy (anonymous):

\[f(x,y)=e ^{\frac{ x }{ y }}\]

OpenStudy (anonymous):

you are right in doing partial differentiation.

OpenStudy (anonymous):

because \[\frac{ d }{ dx }e ^{x}=(1)e^x=e^x\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

@abhrag321 another question though: wouldn't the partial of (x/y) need the quotient rule to be differentiated. so it would be\[\frac{ \partial }{ \partial x }(\frac{ x }{ y })=\frac{ 1(y)-x(0) }{ y^2 }=\frac{ y }{ y^2 }=\frac{ 1 }{ y }\]

OpenStudy (anonymous):

lol my bad im trippin HARD!

OpenStudy (anonymous):

thats the same answer

OpenStudy (anonymous):

yes